OpenStudy (anonymous):

rank stands 450 feet from the base of the Statue of Liberty. If the Statue of Liberty is 306 feet tall, what is the angle of depression from the top of the statue’s torch to Frank? Round the answer to the nearest tenth.

pooja195 (pooja195):

We have a right angled triangle with an opposite of 300.5 ft. (306 - 5.5) and an adjacent of 400 ft. Recalling SOH CAH TOA, tanθ = O/A. tan(θ) = 300.5/400. θ = tan^-1(300.5/400). θ = 36.9°.

pooja195 (pooja195):

\(\huge\color{orange}{WELCOME~TO~OPENSTUDY!!!}\) Refresh your page